Two spherical objects have masses of 3.1 x 105 kg and 6.5 x 103 kg. The gravitational attraction between them is 65 N. How far apart are their centers?


1
Expert's answer
2022-02-15T09:58:06-0500

Given:

"m_1=3.1*10^5\:\rm kg"

"m_2=6.5*10^3\:\rm kg"

"F=65\:\rm N"


The law of universal gravitation says

"F=G\frac{m_1m_2}{r^2}"

Thus

"r=\sqrt{G\frac{m_1m_2}{F}}\\
=\sqrt{6.67*10^{-11}\frac{3.1*10^5*6.5*10^3}{65}}=0.045\:\rm m"